John has a salary of $99,999 and he pays 15,000 in taxes. What percentage does he pay in taxes?
Answer by stanbon(57399)

John has a salary of $99,999 and he pays 15,000 in taxes. What percentage does he pay in taxes?
-----
Proportion:
x/100 = 15000/99999
---
x = 100(0.1500)
x = 15%
==============
Cheers,
Stan H.
==============